the sum of two consecutive multiples of 17 is 85. find the number
[i know the answer is 34 and 51, but how come?

Answer by richwmiller(17219)

17x+17(x+1)=85
17*(2x+1)=85
2x+1=5
2x=4
x=2
17*2+17*3=85
34+51=85